# Continuity of extensions of Lipschitz maps and of monotone maps

### v1

@@ Abstract.
Morever, if either $m\in\{1,2,3\}$ or $X$ is convex, we prove the converse: we show that a map $v\colon X\to\mathbb{R}^{m}$ that allows for a $1$-Lipschitz, uniform distance preserving extension of any $1$-Lipschitz map on a subset of $X$ also satisfies the above set of inequalities.
@@ 1. Introduction
A theorem of Kirszbraun, see [26], tells that any $1$-Lipschitz map on a subset of a finite-dimensional Hilbert space may be extended to a $1$-Lipschitz map on the entire space. It has been later generalised to $1$-Lipschitz maps between arbitrary Hilbert spaces by Valentine, see [42].
@@
Phrased in words, we are interested in whether it is possible to find a $1$-Lipschitz extension of $u$ to $X$ that would be not further to $v$ on $X$ than $u$ is to $v$ on $A$. The main result that we shall prove is the following theorem.
@@
    for all finite $m\leq\mathrm{dim}Y$, all $x,x_{1},\dotsc,x_{m}\in X$ and all non-negative real numbers $t_{1},\dotsc,t_{m}$ that sum up to one there is
@@
    $v(x)-u(x)\in K$
    there exists $1$-Lipschitz extension $\tilde{u}\colon X\to Y$ of $u$ such that for all $x\in X$
    $v(x)-\tilde{u}(x)\in K,$
@@
    $\lVert v(x)-u(x)\rVert\leq\delta$
    there exists $1$-Lipschitz extension $\tilde{u}\colon X\to Y$ of $u$ such that for all $x\in X$
    $\lVert v(x)-\tilde{u}(x)\rVert\leq\delta.$
@@
Let us observe a relation of Theorem 1.2 to the fundamental theorem of affine geometry, see e.g. [4, Theorem 5.4.]. The theorem says that an injective map $v\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$, with $n\geq 2$, that maps any three collinear points onto collinear points, is necessarily affine. As we shall see in Proposition 4.11, if $X$ is convex, and if $\mathrm{dim}Y\geq 2$, then any map satisfying i), is affine. Thus, Theorem 1.2 gives a way to characterise affine and $1$-Lipschitz maps with values in at least two-dimensional space, that resembles the fundamental theorem of affine geometry. Let us note that both characterisations of affine maps do not hold in the one-dimensional setting, see [16, Proposition 3.8].
@@ 1.1. Motivation – multi-dimensional localisation
The motivation for this work stems from considerations pertaining to variational problems concerning $1$-Lipschitz maps, including in particular the optimal transport problem, see [43, 44] for an extensive account on this topic. As described in detail in [16], when the cost function in an optimal transport problem is a metric, then the existence of continuous $1$-Lipschitz extensions of real-valued functions can be used to prove the so-called mass balance condition, which plays a crucial rôle in the localisation technique. The technique was developed in the setting of Riemannian manifolds by Klartag in [27], and in the setting of metric measure spaces by Cavalletti and Mondino in [12, 13], and allowed for proofs of various novel and sharp geometric and functional inequalities. Recent works [19, 20] showed a relation of the localisation technique to martingale transport.
@@
where $\mu$ is a $\mathbb{R}^{m}$-valued Borel measure such that $\mu(\mathbb{R}^{n})=0$, then a multi-dimensional mass balance condition will also be satisfied. That is, let $v\colon\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a $1$-Lipschitz map that attains the above supremum. A leaf $\mathcal{S}$ of $v$ is a maximal subset of $\mathbb{R}^{n}$ such that $v|_{\mathcal{S}}$ is an isometry. The conjecture says if $\mu$ is absolutely continuous, then $\mu(A)=0$ for any Borel set $A$ that is a union of leaves of $v$. As shown in [16, Theorem 3.9], the argument outlined in [27] contains a gap. In [18, Theorem 5] it is shown that in fact the conjecture is false if $m>1$. Theorem 1.2 shows that if $\mu$ is concentrated on a set $X$, and if the maximiser satisfies i) of Theorem 1.2, then the conjecture of Klartag holds true. We shall not describe in detail the proof. Instead, we point out that the proof presented in [27, Corollary 4.5.] works, mutatis mutandis, in this case. Let us point out work [17], where the partitioning of $\mathbb{R}^{n}$ into leaves of a given $1$-Lipschitz map is studied thoroughly.
@@ 1.2. Kirszbraun theorem and related literature
Among a number of the proofs of the Kirszbraun theorem some, e.g. those presented in [26, 42, 41], use the Kuratowski–Zorn lemma and thus are non-constructive. Other proofs, among them [2, 10], are constructive. There exists also a formula for the extension, see [5]. Let us also note approaches that use Fenchel duality and Fitzpatrick functions, see [7, 40]. Related methods also allowed for a constructive proof, see [8]. These methods allow for further and novel contributions to the field. Let us mention a recent work [11], where an invariant version of the Kirszbraun theorem is proven.
We refer the reader also to work of Dacorogna and Gangbo [22] where various extensions properties of vector-valued maps are studied. In [14] a version of the Kirszbraun theorem for graphs is provided.
@@ 1.3. Continuity of extensions
In [16] we studied the rate of continuity of extensions of Lipschitz maps. Namely, we supposed that we are given two sets $A\subset B\subset\mathbb{R}^{n}$ and $1$-Lipschitz maps $u\colon A\to\mathbb{R}^{m}$ and $v\colon B\to\mathbb{R}^{m}$, with $m>1$. We were interested in
@@
In [28, 29, 30] a related question of the existence of a Lipschitz constant preserving continuous extension operator on the space of Lipschitz maps is studied.
@@ 1.4. Monotone maps
Existence of maximal monotone operators is well-established to be related to extensions of Lipschitz maps. The topic of maximal monotone operators has been initiated by Minty, [35, 36], studied in [23] and has since been an active area of research. Let us mention recent works [24, 11, 45] pertaining to the topic. However, to the author knowledge, no results concerining continuity of monotone extensions are known.
Let $v\colon X\to Y^{*}$ and let $A\subset X$. We shall study conditions on $v$ under which for any monotone $u\colon A\to Y^{*}$, there exists a monotone extension $\tilde{u}\colon X\to Y^{*}$ of $u$ that is no further to $v$ on $X$ than $u$ is to $v$ on $A$.
@@
Let $Y$ be a linear topological space. Suppose that $X\subset Y$. Let $v\colon X\to Y^{*}$. Suppose that for all finite $m\leq\mathrm{dim}(Y^{*})$, all $x,x_{1},\dotsc,x_{m}\in X$ and all non-negative real numbers $t_{1},\dotsc,t_{m}$ that sum up to one there is
$\Big{\langle}v(x)-\sum_{i=1}^{m}t_{i}v(x_{i}),x-\sum_{i=1}^{m}t_{i}x_{i}\Big{\rangle}\geq 0,$
@@ 1.5. Maps of semi-bounded strain
In the study of Michell trusses, see e.g. [32, 33, 34] for foundational works and and [6, 9, 38, 39] for recent contributions, a class of functions of bounded strain plays an important rôle. The class consists of functions $u\colon A\to\mathbb{R}^{n}$, where $A\subset\mathbb{R}^{n}$, such that
@@
The extensions properties of this class has been studied by Dacorogna and Gangbo in [22]. In particular, it is shown in [22, Theorem 22.] that there exist sets $A\subset\mathbb{R}^{n}$ and functions $u\colon A\to\mathbb{R}^{n}$ satisfying (1.1) with no extension to $\mathbb{R}^{n}$.
@@
In [22, Proposition 28., (i)], it is shown that if $A$ is a finite set, then $u$ can be extended to a larger finite set. However, as shown in [22, Proposition 28., (ii)], there are infinite sets $A$ and functions $u$ satisfying (1.2), with no extension to a larger set.
@@
Suppose now that $v\colon X\to\mathbb{R}^{n}$, $A\subset X$. Here we that, under certain conditions on $v$, not only does there exists an extension of any $u\colon A\to\mathbb{R}^{n}$ of $1$-semi-bounded strain to a map $\tilde{u}\colon X\to\mathbb{R}^{n}$ of $1$-semi-bounded strain on $X$, but also such $\tilde{u}$ can be chosen to be not further to $v$ on $X$ than $u$ is to $v$ on $A$.
@@
Let $Y$ be a real Hilbert space. Suppose that $X\subset Y$. Let $v\colon X\to Y$. Suppose that for all finite $m\leq\mathrm{dim}Y$, all $x,x_{1},\dotsc,x_{m}\in X$ and all non-negative real numbers $t_{1},\dotsc,t_{m}$ that sum up to one there is
@@
The theorem is, to the author knowledge, the first result that gives a condition under which a map of $1$-semi-bounded strain admits and extension and the first result concerned with continuity of extensions of maps of $1$-semi-bounded strain.
@@ 2. Minty theorem
We shall start by recalling one of the main tools that we shall use in the proof of Theorem 1.2. It is the theorem of Minty, see [37], which encompasses several Kirszbraun-type theorems. We shall first bring to the reader attention the following definition.
@@
Let $Y$ be a real vector space and $X$ be a set. A real function on $Y$ is called finitely lower semi-continuous if its restriction to any finitely-dimensional subspace of $Y$ is lower semi-continuous. A function $\Phi\colon Y\times X\times X\to\mathbb{R}$ is called a $K$-function if $\Phi$ is both finitely lower semi-continuous and convex in the first variable and such that for any points $(y_{i},x_{i})_{i=1}^{l}\in Y\times X$, $x\in X$ and any $\lambda_{1},\dotsc,\lambda_{l}\geq 0$ such that $\sum_{i=1}^{l}\lambda_{i}=1$ we have
@@
Let $Y$ be a real vector space and $X$ be a set. Let $\Phi\colon Y\times X\times X\to\mathbb{R}$ be a $K$-function. Let
@@ 3. Existence of extensions of $1$-Lipschitz maps
$\Phi(y,x,x^{\prime})=\lVert y\rVert^{2}+2\langle y,v(x)-v(x^{\prime})\rangle-\lVert x-x^{\prime}\rVert^{2}+\lVert v(x)-v(x^{\prime})\rVert^{2},y\in V,x,x^{\prime}\in X.$
Let us check that it is a $K$-function, see Definition 2.1. The condition of convexity and finitely lower semi-continuity is clearly satisfied. We need only to check whether the condition (2.1) holds, with $l\leq m$, as $\mathrm{dim}V\leq k-1\leq m-1$. It is readily seen that the first two summands in the definition of $\Phi$ both satisfy the condition (2.1) with equalities. Indeed, let $l\leq m$ and $(y_{i},x_{i})_{i=1}^{l}\in Y\times X$, $x\in X$ and let $\lambda_{1},\dotsc,\lambda_{l}\geq 0$ be such that $\sum_{i=1}^{l}\lambda_{i}=1$. Then a short calculation readily implies that
@@
is non-empty. Indeed, condition that $\Phi(y-y_{i},x_{i},x)\leq 0$, for $i=1,\dotsc,k$ and $y_{i}=u(x_{i})-v(x_{i})$, is equivalent to that for $i=1,\dotsc,k$
$\lVert y-v(x)-(y_{i}-v(x_{i}))\rVert\leq\lVert x-x_{i}\rVert.$
@@
Note now that the balls are compact in $Y$, thanks to the Banach–Alaoglu theorem. Therefore for $x\in X$, the intersection of the form
$\bigcap\Big{\{}B(u(x^{\prime}),\lVert x-x^{\prime}\rVert)\cap(v(x^{\prime})+K)\mid x^{\prime}\in A\Big{\}}$
@@ 4.1. Leaves of $1$-Lipschitz maps
We shall recall [17, Lemma 2.4.]. Although it is formulated in [17] merely for finite-dimensional spaces, the proof works mutatis mutandis also for general Hilbert spaces.
Let $A\subset Z$. We shall say that a map $v\colon A\to Y$ is an isometry provided that for all $x,y\in A$ there is $\lVert v(x)-v(y)\rVert=\lVert x-y\rVert$.
@@
We cite the following definition, see [17, Definition 2.2.], that is motivated by the above lemma. We refer the reader to [17] for a detailed study of this notion.
@@
Then for all $x_{1},\dotsc,x_{m}\in A$ and all non-negative $t_{1},\dotsc,t_{m}$ that sum up to one there is

### v2

@@ Abstract.
Moreover, if either $m\in\{1,2,3\}$ or $X$ is convex, we prove the converse: we show that a map $v\colon X\to\mathbb{R}^{m}$ that allows for a $1$-Lipschitz, uniform distance preserving extension of any $1$-Lipschitz map on a subset of $X$ also satisfies the above set of inequalities.
@@ 1. Introduction
A theorem of Kirszbraun, see [26], tells that any $1$-Lipschitz map on a subset of a finite-dimensional Hilbert space may be extended to a $1$-Lipschitz map on the entire space. This result has been later generalised by Valentine, see [42], so that it is applicable to $1$-Lipschitz maps between arbitrary Hilbert spaces
@@
Phrased in words, we are interested in whether it is possible to find a $1$-Lipschitz extension of $u$ to $X$ that would be not farther to $v$ on $X$ than $u$ is to $v$ on $A$. The main result that we shall prove is the following theorem.
@@
    for all finite $m\leq\mathrm{dim}Y$, all $x,x_{1},\dotsc,x_{m}\in X$ and all non-negative real numbers $t_{1},\dotsc,t_{m}$ that sum up to one we have
@@
    $u(x)-v(x)\in K$
    there exists a $1$-Lipschitz extension $\tilde{u}\colon X\to Y$ of $u$ such that for all $x\in X$
    $\tilde{u}(x)-v(x)\in K,$
@@
    $\lVert u(x)-v(x)\rVert\leq\delta$
    there exists a $1$-Lipschitz extension $\tilde{u}\colon X\to Y$ of $u$ such that for all $x\in X$
    $\lVert\tilde{u}(x)-v(x)\rVert\leq\delta.$
@@
Let us observe a relation of Theorem 1.2 to the fundamental theorem of affine geometry, see e.g., [4, Theorem 5.4.]. The theorem says that an injective map $v\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$, with $n\geq 2$, that maps any three collinear points onto collinear points, is necessarily affine. Clearly, if $X$ is convex, and if $\mathrm{dim}Y\geq 2$, then any map satisfying i) is affine. Thus, Theorem 1.2 gives a way to characterise affine and $1$-Lipschitz maps with values in an at least two-dimensional space, that resembles the fundamental theorem of affine geometry. Let us note that both characterisations of affine maps do not hold in the one-dimensional setting, see [16, Proposition 3.8].
@@ 1.1. Motivation – multi-dimensional localisation
The motivation for this work stems from considerations pertaining to variational problems concerning $1$-Lipschitz maps, including in particular the optimal transport problem, see [43, 44] for an extensive account on this topic. As described in detail in [16], when the cost function in an optimal transport problem is a metric, then the existence of continuous $1$-Lipschitz extensions of real-valued functions can be used to prove the so-called mass balance condition, which plays a crucial role in the localisation technique. The technique was developed in the setting of Riemannian manifolds by Klartag in [27], and in the setting of metric measure spaces by Cavalletti and Mondino in [12, 13], and allowed for proofs of various novel and sharp geometric and functional inequalities. Recent works [19, 20] showed a relation of the localisation technique to martingale transport.
@@
where $\mu$ is a $\mathbb{R}^{m}$-valued Borel measure such that $\mu(\mathbb{R}^{n})=0$, then a multi-dimensional mass balance condition will also be satisfied. That is, let $v\colon\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a $1$-Lipschitz map that attains the above supremum. A leaf $\mathcal{S}$ of $v$ is a maximal subset of $\mathbb{R}^{n}$ such that $v|_{\mathcal{S}}$ is an isometry. The conjecture says if $\mu$ is absolutely continuous, then $\mu(A)=0$ for any Borel set $A$ that is a union of leaves of $v$. As shown in [16, Theorem 3.9], the argument outlined in [27] contains a gap. In [18, Theorem 5] it is shown that in fact the conjecture is false if $m>1$. Theorem 1.2 shows that if $\mu$ is concentrated on a set $X$, and if the maximiser satisfies i) of Theorem 1.2, then the conjecture of Klartag holds true. We shall not describe in detail the proof. Instead, we point out that the proof presented in [27, Corollary 4.5] works, mutatis mutandis, in this case. Let us point out work [17], where the partitioning of $\mathbb{R}^{n}$ into leaves of a given $1$-Lipschitz map is studied thoroughly.
@@ 1.2. Kirszbraun theorem and related literature
Among a number of the proofs of the Kirszbraun theorem some, e.g., those presented in [26, 42, 41], use the Kuratowski–Zorn lemma and thus are non-constructive. Other proofs, among them [2, 10], are constructive. There exists also a formula for the extension, see [5]. Let us also note the approaches that use the Fenchel duality and Fitzpatrick functions, see [7, 40]. Related methods also allowed for a constructive proof, see [8]. These methods allow for further and novel contributions to the field. Let us mention a recent work [11], where an invariant version of the Kirszbraun theorem is proven.
We refer the reader also to the work of Dacorogna and Gangbo [22] where various extensions properties of vector-valued maps are studied. In [14] a version of the Kirszbraun theorem for graphs is provided.
@@ 1.3. Continuity of extensions
In [16] we studied the rate of continuity of extensions of Lipschitz maps. Namely, we supposed that we were given two sets $A\subset B\subset\mathbb{R}^{n}$ and $1$-Lipschitz maps $u\colon A\to\mathbb{R}^{m}$ and $v\colon B\to\mathbb{R}^{m}$, with $m>1$. We were interested in
@@
In [28, 29, 30] a related question concerning the existence of a Lipschitz constant preserving continuous extension operator on the space of Lipschitz maps is studied.
@@ 1.4. Monotone maps
Existence of maximal monotone operators is well-established to be related to extensions of Lipschitz maps. The topic of maximal monotone operators has been initiated by Minty, [35, 36], studied in [23], and has since been an active area of research. Let us mention recent works [11, 24, 45] pertaining to the topic. However, to the author’s knowledge, no results concerning continuity of monotone extensions are known.
Let $v\colon X\to Y^{*}$ and let $A\subset X$. We shall study conditions on $v$ under which for any monotone $u\colon A\to Y^{*}$, there exists a monotone extension $\tilde{u}\colon X\to Y^{*}$ of $u$ that is no farther to $v$ on $X$ than $u$ is to $v$ on $A$.
@@
Let $Y$ be a linear topological space. Suppose that $X\subset Y$. Let $v\colon X\to Y^{*}$. Suppose that for all finite $m\leq\mathrm{dim}(Y^{*})$, all $x,x_{1},\dotsc,x_{m}\in X$ and all non-negative real numbers $t_{1},\dotsc,t_{m}$ that sum up to one we have
$\Big{\langle}v(x)-\sum_{i=1}^{m}t_{i}v(x_{i}),x-\sum_{i=1}^{m}t_{i}x_{i}\Big{\rangle}\geq 0.$
@@ 1.5. Maps of semi-bounded strain
In the study of Michell trusses, see, e.g., [32, 33, 34] for foundational works and and [6, 9, 38, 39] for recent contributions, a class of functions of bounded strain plays an important role. The class consists of functions $u\colon A\to\mathbb{R}^{n}$, where $A\subset\mathbb{R}^{n}$, such that
@@
The extensions properties of this class have been studied by Dacorogna and Gangbo in [22]. In particular, it is shown in [22, Theorem 22] that there exist sets $A\subset\mathbb{R}^{n}$ and functions $u\colon A\to\mathbb{R}^{n}$ satisfying (1.1) with no extension to $\mathbb{R}^{n}$.
@@
In [22, Proposition 28, (i)], it is shown that if $A$ is a finite set, then $u$ can be extended to a larger finite set. However, as shown in [22, Proposition 28, (ii)], there are infinite sets $A$ and functions $u$ satisfying (1.2), with no extension to a larger set.
@@
Suppose now that $v\colon X\to\mathbb{R}^{n}$, $A\subset X$. Here we show that, under certain conditions on $v$, not only does there exist an extension of any $u\colon A\to\mathbb{R}^{n}$ of $1$-semi-bounded strain to a map $\tilde{u}\colon X\to\mathbb{R}^{n}$ of $1$-semi-bounded strain on $X$, but also such $\tilde{u}$ can be chosen to be no farther to $v$ on $X$ than $u$ is to $v$ on $A$.
@@
Let $Y$ be a real Hilbert space. Suppose that $X\subset Y$. Let $v\colon X\to Y$. Suppose that for all finite $m\leq\mathrm{dim}Y$, all $x,x_{1},\dotsc,x_{m}\in X$ and all non-negative real numbers $t_{1},\dotsc,t_{m}$ that sum up to one we have
@@
The theorem is, to the author’s knowledge, the first result that gives a condition under which a map of $1$-semi-bounded strain admits an extension and the first result concerned with continuity of extensions of maps of $1$-semi-bounded strain.
@@ 2. Minty theorem
We shall start by recalling one of the main tools that we shall use in the proof of Theorem 1.2. It is the theorem of Minty, see [37], which encompasses several Kirszbraun-type theorems. We shall first bring to the reader’s attention the following definition.
@@
Let $Y$ be a real vector space and $X$ be a set. A real function on $Y$ is called finitely lower semi-continuous if its restriction to any finitely-dimensional subspace of $Y$ is lower semi-continuous. A function $\Phi\colon Y\times X\times X\to\mathbb{R}$ is called a Kirszbraun function if $\Phi$ is both finitely lower semi-continuous and convex in the first variable and for any natural number $l$, any points $(y_{i},x_{i})_{i=1}^{l}\in Y\times X$, $x\in X$ and any $\lambda_{1},\dotsc,\lambda_{l}\geq 0$ such that $\sum_{i=1}^{l}\lambda_{i}=1$ we have
@@
Definition 2.2.
For a subset $S\subset Y$ of a real vector space $Y$ we denote by $\mathrm{Conv}S$ the smallest convex set in $Y$ containing $S$. It shall be called the convex hull of $S$.
@@
Let $Y$ be a real vector space and $X$ be a set. Let $\Phi\colon Y\times X\times X\to\mathbb{R}$ be a Kirszbraun function. Let
@@ 3. Existence of extensions of $1$-Lipschitz maps
$\Phi(y,x,x^{\prime})=\lVert y\rVert^{2}-2\langle y,v(x)-v(x^{\prime})\rangle-\lVert x-x^{\prime}\rVert^{2}+\lVert v(x)-v(x^{\prime})\rVert^{2},y\in V,x,x^{\prime}\in X.$
Let us check that it is a Kirszbraun function, see Definition 2.1. The condition of convexity and finitely lower semi-continuity is clearly satisfied. We need only to check whether the condition (2.1) holds, with $l\leq m$, as $\mathrm{dim}V\leq k-1\leq m-1$. It is readily seen that the first two summands in the definition of $\Phi$ both satisfy the condition (2.1) with equalities. Indeed, let $l\leq m$ and $(y_{i},x_{i})_{i=1}^{l}\in V\times X$, $x\in X$ and let $\lambda_{1},\dotsc,\lambda_{l}\geq 0$ be such that $\sum_{i=1}^{l}\lambda_{i}=1$. Then a short calculation readily implies that
@@
is non-empty. Indeed, let us put $y_{i}=u(x_{i})-v(x_{i})$ for $i=1,2,\dotsc,k$. Then
$\Phi(y_{i}-y_{j},x_{i},x_{j})\leq 0\text{ for }i,j=1,2,\dotsc,k$
is equivalent to $u\colon\{x_{1},\dotsc,x_{k}\}\to Y$ being $1$-Lipschitz. Now, Theorem 2.3 ensures that there exists $y\in Y$ such that
$\Phi(y-y_{i},x_{i},x)\leq 0\text{, for }i=1,\dotsc,k,$
and such that $y\in K$, since $y\in\mathrm{Conv}\{y_{1},\dotsc,y_{k}\}$. The definition of $\Phi$ shows that this $y$ satisfies
$\lVert y+v(x)-(y_{i}+v(x_{i}))\rVert\leq\lVert x-x_{i}\rVert\text{ for }i=1,2,\dotsc,k.$
@@
Thanks to the Banach–Alaoglu theorem the balls in $Y$ are compact in the weak* topology. Therefore for $x\in X$, the intersection of the form
$\bigcap\Big{\{}B(u(x^{\prime}),\lVert x-x^{\prime}\rVert)\cap(v(x)+K)\mid x^{\prime}\in A\Big{\}}$
@@ 4.1. Leaves of $1$-Lipschitz maps
We shall recall [17, Lemma 2.4]. Although it is formulated in [17] merely for finite-dimensional spaces, the proof works mutatis mutandis also for general Hilbert spaces.
Let $A\subset Z$. We shall say that a map $v\colon A\to Y$ is an isometry provided that for all $x,y\in A$ we have $\lVert v(x)-v(y)\rVert=\lVert x-y\rVert$.
@@
We cite the following definition, see [17, Definition 2.2], that is motivated by the above lemma. We refer the reader to [17] for a detailed study of this notion.
@@
Then for all $x_{1},\dotsc,x_{m}\in A$ and all non-negative $t_{1},\dotsc,t_{m}$ that sum up to one we have